OpenStudy (anonymous):

Help? I'll post the Q down below(:

OpenStudy (anonymous):

Chose the equivalent value \[5\sqrt{3}\] (A) \[\sqrt{15}\] (B) \[\sqrt{53}\] (C) \[\sqrt{75}\] (D) 225 I don't know what I should do.

OpenStudy (mathmate):

Hint: use the identity a*sqrt(b)=sqrt(a^2*b) and check the different answers.

OpenStudy (nory):

Did you learn about simplifying radicals? For example, √200 = √2 * √100 = 10√2? If so, try simplifying A, B, C, and D and seeing what you get. If not, tell me and I can explain it in more detail.

OpenStudy (anonymous):

ohh..Okay! And yeah. I learned how to simplify radicals.(: Thanks.

OpenStudy (nory):

No problem. :) Thanks for the medal.